OpenStudy (anonymous):

whats the percent of decrease... 18 to 10?? how do i do these?? please and thank you

OpenStudy (anonymous):

...Nightie's got the conch.

OpenStudy (anonymous):

Nah, it's ok, I'm going to uni now :P

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

magic conch can i have a burger?

OpenStudy (anonymous):

In these types of questions, you're always looking for something called 'relative difference' and it's *defined* like this:\[\frac{(final)-(initial)}{(initial)}\]where 'final' and 'initial' are the final and initial values respectively. You have\[\frac{10-18}{18}=-\frac{4}{9} =-0.4444....\]To get percent, you multiply it by 100%;\[-\frac{4}{9} \times 100 \approx -44.44\]percent.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

no worries.